OpenStudy (anonymous):

Medal! Which ordered pair is the solution to the system of equations? 2x+y=-8 x-y=-4 A) (-3,-2) B) (-1,3) C) (1,5) D) (-4,0)

OpenStudy (anonymous):

ok so i would say D because 2*-4+0=-8 -8+0=-8 -8=-8 ok so it works for that one now the second one -4-0=-4- -4=-4